Section 3: (10 marks) Word Problem

Target response size: text portion approximately 1 - 2 paragraphs (or equivalent) of explanation, plus maths portion showing your work

Four electrical devices are connected to powerpoints wired into the same electrical supply circuit. This is a standard Australian residential building circuit rated at 10 A. The devices have power ratings of 200 W, 740 W, 330 W and 1150 W. Can all four devices be operated at the same time on this circuit? 


1
Expert's answer
2021-05-04T06:16:18-0400

We have given the current rating i.e. 10A

We know Voltage rating in Australia = 230V

For 1st Device,

"P=VI"

"I_1 = \dfrac{200}{230} = 0.86A"

Similarly for the other three devices current will be

"I_2 = 3.21A"

"I_3=1.43A"

"I_4=5A"

Total current "= 5+1.43+3.21+0.86 = 10.5A"

Hence this value ids greater than the Rated value i.e. "10A."

Hence, all four devices cannot be operated at the same time on this circuit.